QUESTION: 1 (6 marks) In March 2019 AAR Ltd reported net profits after tax of $500,000 for the year 2018 andannounced that it will pay the annual dividend on 1 June 2019. AAR expects the net profits after tax for 2019 to be 20% higher and it will maintain its current dividend payout ratio of 60%. AAR will pay the annual dividend for 2019 in 1 June, 2020. Karina Adams owns 10% of the ordinary share capital of AAR Pty. She has no savings and her sole income comes from AAR Pty. If Karina wishes to buy $5000 worth of furniture in June 2019 and pays all her other bills totaling $35,000, calculatehow much can she consume June 2020? Assume the market rate of interest is 8% per annum
QUESTION 2 (3+58 marks) James King bought a house three years ago that cost $750,000. James put up 20% deposit and borrowed the rest from FC Bank at a rate of 7.2% per annum, compounded monthly, for 10 years. Three months ago FC Bank notified James that after the last monthly payment for the third year, the interest rate on his loan will increase to 9.6% per annum, compounded monthly, in line with market rates. Also from the fourth year of his loan James can either increase the monthly repayment (so as to pay off the loan by the originally agreed date), or he can keep paying the same original monthly repayment and extend the term of the loan. Calculate the new monthly repayment if Jamespay off the loan by the originally a. agreed date b. Calculate if James keeps on paying the original repayment the extra period added to the term of the loan
Answer #1

Solution :1 Dividend= 500000*60%=$300000

Karina's Share= $300000*10%=$30000

Expenses = $5000=$35000=40,000

Net amount to be be borrowed= $10,000

On June 2020: N.P increase by 20% i.e $600000

Dividend Distributed: $600000860%=$360000

Karina's Share= $360000*10%=$36000

Interest Amount: 10,000@8%=$800

Amount payable= $10,000+800=$10,800

Total Karina cam consume in June 2020= $25200 (36000-10800).

Solution 2: a) New EMI shall be $7586

b) Tenure shall increase by 6 months